OpenStudy (anonymous):

((b^n)^2)/((b^n)*(b^n+2)) Write as a power of b. ((1)-(y^-1))/((y)-(y^-1)) Simplify. (2^x) + (2^-x) = (5/2) Solve for x. ((b^n)^2)/((b^n)*(b^n+2)) Write as a power of b. ((1)-(y^-1))/((y)-(y^-1)) Simplify. (2^x) + (2^-x) = (5/2) Solve for x. @Mathematics

OpenStudy (jamesj):

((b^n)^2)/((b^n)*(b^n+2)) \[ = \frac{(b^n)^2}{b^n.b^{n+2}} = \frac{b^{2n}}{b^{2n+2}} = b^{2n}.b^{-(2n+2)} = b^{2n - 2n - 2} = b^{-2} \]

OpenStudy (jamesj):

= 1/b^2

OpenStudy (jamesj):

For the second equation, multiply top and bottom by y

OpenStudy (jamesj):

For the third equation: (2^x) + (2^-x) = (5/2) 1. Multiply both sides by 2^x 2. Set u = 2^x 3. Turn the resulting equation from step 1 into a quadratic equation in u 4. Solve that equation for u 5. Now set u back equal to 2^x and solve for x.

OpenStudy (jamesj):

...for second equation, after multiplying top and bottom by y, then factorize and simplify.

OpenStudy (anonymous):

what's y times (1-(y^-1)) and what's y times (y-(y^-1))

OpenStudy (jamesj):

\[ y(1-y^{-1}) = y(1 - 1/y) = y - (y/y) = y - 1\] You figure out the other one.

OpenStudy (anonymous):

after I find both, then what?

OpenStudy (anonymous):

nvm I got it

OpenStudy (anonymous):

how about last one

OpenStudy (jamesj):

Look at what you've got and see if you can simplify further, which you should be able to. For the last one, I wrote down the approach for you above in steps.

OpenStudy (anonymous):

I don't get how u do it

OpenStudy (jamesj):

For the third equation: (2^x) + (2^-x) = (5/2) 1. Multiply both sides by 2^x 2. Set u = 2^x 3. Turn the resulting equation from step 1 into a quadratic equation in u 4. Solve that equation for u 5. Now set u back equal to 2^x and solve for x. ====== 1. Multiply both sides by 2^x \[ (2^x)^2 + 2^{-x}.2^x = 5/2 . 2^x \] \[ (2^x)^2 + 1 = 5/2 . 2^x \] 2. Set u = 2^x \[ u^2 + 1 = 5/2 . u \] 3. Turn the resulting equation from step 1 into a quadratic equation in u \[ u^2 - (5/2)u + 1 = 0 \] You take it from here.

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

what's the period mean?

OpenStudy (anonymous):

after the 2s?

OpenStudy (jamesj):

multiplication

OpenStudy (anonymous):

I can't factor it :(

OpenStudy (anonymous):

and btw, the answers are -1 and 1, but idk how